### v1

# Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations

Abstract
Large language models like GPT, LLAMA, and Claude have become incredibly powerful at generating text, but they are still black boxes, so it is hard to understand how they decide what to say. That lack of transparency can be problematic, especially in fields where trust and accountability matter. To help with this, we introduce SMILE, a new method that explains how these models respond to different parts of a prompt. SMILE is model-agnostic and works by slightly changing the input, measuring how the output changes, and then highlighting which words had the most impact. Create simple visual heat maps showing which parts of a prompt matter the most. We tested SMILE on several leading LLMs and used metrics such as accuracy, consistency, stability, and fidelity to show that it gives clear and reliable explanations. By making these models easier to understand, SMILE brings us one step closer to making AI more transparent and trustworthy.
Figure 1: Highlighting the contribution of each word in the input prompt of a Commercial LLM
1 Introduction
Large Language Models (LLMs) have become powerful tools in recent years. They can perform tasks ranging from answering questions with remarkable accuracy [1] to drafting essays and other written content with coherence and style [2], creating art that spans various mediums and styles [3], coding across multiple programming languages [4], debugging software [4], writing technical documentation [5], and even generating algorithms [1].
LLMs function by leveraging deep learning architectures, particularly the Transformer model. This model uses self-attention mechanisms to process and generate text by analysing relationships between words in large data sets [6]. These models are trained in various corpora, including books, articles, code, and web content, allowing them to learn language patterns, context, and syntax on an unheard scale [7].
Their capabilities continue to grow, turning them into essential tools for a wide range of tasks [2]. These models, trained in various texts, have transformed how we interact with technology and how machines understand and generate human language [6]. LLMs such as GPT, BERT, and Claude leverage the Transformer architecture to process and generate language based on patterns learned from extensive datasets [7]. Although they excel in many NLP tasks, understanding their strengths and limitations is the key to using them effectively and responsibly [8].
These models continue to evolve, pushing the boundaries of AI applications in language understanding and generation [2]. However, the nature of these models is known as black boxes and lacks transparency [9]. One of their important problems is how these models decide which words or parts of input to focus on, which remains a complex and fascinating process [10].
At their core, LLMs rely on patterns and relationships learned during training to decide which input elements are the most important [6]. This process is more than just identifying individual keywords; it is about picking up sophisticated connections and context that shape the model’s responses [10]. Understanding how this works is crucial: it explains why these models excel in some scenarios but struggle in others. It also highlights why they can occasionally produce responses that are biased, irrelevant, or even harmful [11].
By studying how LLMs emphasise words and phrases, we can better understand how they work [10]. This can help us make them more transparent and reliable, ensuring they focus on what truly matters in a given context. It also allows us to address potential problems, such as biases or alignment problems, making these tools safer and more ethical [11].
As LLMs continue to integrate into our daily lives and decision-making processes, understanding their inner workings is more important than ever [9]. Research on how they prioritise language helps us refine their performance, ensure ethical use, and unlock their full potential to benefit society [11] responsibly.
To solve this problem, we propose a novel approach of SMILE (Statistical Model-agnostic Interpretability with Local Explanations) to extract a visual heat map that shows the influence of each word in responding to these models [12]. This method extends the foundation laid by LIME [13] by incorporating Empirical Cumulative Distribution Function (ECDF) statistical distances, significantly improving robustness and making interpretability more effective.
SMILE produces visual heatmaps that reveal the impact of individual words or phrases in text input on the model’s output as illustrated in Fig 1. For example, when a language model processes an instruction or query, the heatmap highlights the importance of specific terms in influencing the outcome. This offers users a clear visual representation of how the model interprets and prioritises different parts of the input text [10].
To ensure a thorough evaluation of our approach, we employ a suite of metrics, including accuracy, fidelity, stability, and consistency [14]. These metrics demonstrate that SMILE provides reliable insights into model behaviour [15].
The contributions of our work include the following.
  • Introducing a New Interpretability Framework: SMILE is a cutting-edge method that visualises the role of individual words in shaping model output, offering a deeper understanding of how LLMs process language [12].
  • Promoting Transparency and Trust in LLMs: By visually linking text input to model decisions, SMILE enhances transparency, making LLMs more trustworthy and enabling their practical use in high-stakes domains such as healthcare, education, and legal systems.
  • Comprehensive Evaluation for Robustness: Our evaluation leverages metrics such as stability, fidelity, accuracy, and consistency, establishing a new standard for evaluating the interpretability and reliability of language models [16].
2 Literature review
This section summarises the most recent advances in large language models (LLMs) and their growth in various applications. We also discuss Explainable Artificial Intelligence (XAI), a field dedicated to making AI systems more transparent and easier to understand. Finally, we explore how these two areas come together, focusing on the ongoing efforts to create explainable large language models. We discuss the key challenges involved and the techniques being developed to improve how we interpret and trust these powerful models.
2.1 Large Language Models
Large language models (LLMs) have become increasingly common thanks to significant strides in deep learning, better access to data, and more powerful computing resources. These models are trained on vast amounts of text and can generate responses that sound natural, answer questions, and support a wide range of language-based tasks. Over time, several notable LLMs have helped push the field forward, each bringing new ideas and improvements.
GPT (OpenAI) [1] is one of the most well-known autoregressive transformer-based models, capable of generating coherent and contextually rich text. Google’s T5 [17] introduced the text-to-text framework, where all NLP tasks are framed as text generation problems, improving transfer learning and efficiency. Similarly, PaLM (Google) [18] leverages the Pathways architecture for large-scale language modelling, optimising multitask learning across different NLP tasks.
Anthropic’s Claude [19] focuses on ethical AI and safe interactions designed to minimise harmful outputs while maintaining interpretability. Google’s Gemma 2 [20] is optimised for logical reasoning and structured data interpretation, making it a strong candidate for academic and enterprise use cases. Meta’s LLaMA [5] is an efficient, open-source model tailored for various NLP applications, offering researchers an accessible yet powerful alternative to proprietary models.
EleutherAI’s GPT-NeoX [21] is an open-source alternative to GPT, providing robust text generation capabilities while allowing customisation. AI21 Labs’ Jurassic-2 [22] is designed for large-scale enterprise applications, emphasising contextual understanding and user intent. Cohere’s Command R+ [23] integrates retrieval-augmented generation, improving fact-based responses and minimising hallucinations.
Furthermore, models like BERT (Google) [24] have NLP through bidirectional context understanding, significantly improving sentiment analysis and question-answering tasks. Microsoft’s Phi-4 focuses on efficiency and cost-effectiveness, delivering competitive performance with fewer parameters. Mistral and Alibaba’s Qwen 2.5 push the boundaries of model adaptability and fine-tuning capabilities, making them suitable for multilingual and domain-specific applications.
DeepSeek-LLM [25] is another promising open-source model, competing with top-tier models regarding benchmark performance. Models such as Mixtral and StableLM have been developed with efficiency and specific NLP task performance in mind, ensuring a balance between accuracy and computational demands.
Recent developments have introduced models such as DeepSeek-R1, which focused on advanced reasoning capabilities, and Qwen2.5-Max, which explores large-scale Mixture-of-Experts (MoE) models to enhance efficiency and performance. OpenAI’s o3-mini pushes the frontier of cost-effective reasoning, balancing performance with computational efficiency. DeepSeek-V3, the first open-sourced GPT-4o level model, represents a significant step forward in open-source LLM development.
The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications:
Category Models Description General-Purpose GPT-4 [26], GPT-4o [27], Claude 3 [28], Gemini 1.5 [29], Command R+ [30], Mistral [31] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [32], Mistral [31], Mixtral [33], Falcon 180B [34], DeepSeek-V3 [35], OpenChat [36] Customizable models for research applications Retrieval-Augmented Command R+ [30], GPT-4 Bing [26], Claude 3 RAG [28], Gemini Search [29] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [28], Gemini 1.5 [29], GPT-4o [27], Grok-1.5 [37] Specialized in processing long-form text Advanced Reasoning GPT-4 [26], Claude 3 Opus [28], Gemini 1.5 [29], Mixtral [33], Grok-1.5 [37] Logical reasoning, multi-task learning, problem-solving Ethically-Aligned Claude 3 [28], Gemini [29], GPT-4 [26], Grok [37] Emphasize safety, reducing bias, ethical AI Mixture-of-Experts (MoE) Mixtral [33], Qwen2.5-Max [38], GShard [39], Switch Transformer [40] Improved efficiency and performance via MoE Multimodal GPT-4o [27], Gemini 1.5 [29], Claude 3 [28], Kosmos-2 [41], Grok-1.5V [37] Capable of processing both text and visual inputs Multilingual BLOOM [42], XGLM [43], mBERT [44], XLM-R [45], ByT5 [46] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [47], Galactica [48], Legal-BERT [49], StarCoder2 [50], Code LLaMA [51] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [52], Vicuna [53], ChatGLM [54], Baize [55], OpenChat [36] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [56], TinyLLaMA [57], DistilGPT2 [58], GPT-2 [59], LLaMA 2 7B [60] Designed for fast inference and deployment on edge devices
Figure 2: Extended LLM Categories, Models, and Descriptions
2.2 Explainable AI (XAI)
The increasing complexity of machine learning models has made understanding how they make decisions more important than ever. Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [61, 14].
Explainability methods can be broadly categorised as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [9].
In contrast, post-hoc methods, such as Local Interpretable Model-agnostic Explanations (LIME) [13] and SHAP (Shapley Additive Explanations) [134], are applied after model training to explain the decisions of complex, often “black-box” models like neural networks and ensemble methods. These approaches provide valuable insights without altering the underlying model structure [63]. Post-hoc explainability methods can be classified into global and local explanations.
Global explanations provide an overarching view of the model’s decision-making patterns across an entire dataset, helping to identify trends and biases [62, 64]. In contrast, local explanations focus on specific predictions, providing a detailed breakdown of how individual inputs influence model outputs. LIME and its numerous enhancements fall under local explainability methods and are designed to provide granular, case-specific interpretations [13].
LIME has become one of the most widely used explainability techniques, but over time, researchers have introduced various improvements to address its limitations and expand its usefulness across different applications. As shown in Figure 3, enhancements to LIME can be grouped into four main categories:
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [65], S-LIME [66], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behaviour, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimisation strategies to make LIME more efficient and stable. Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [67], TS-MULE for time series data [68], Graph LIME for graph-structured data [69], Sound LIME for audio data [70], and B-LIME for ECG signal data [71] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to improve LIME’s approach, including using alternative surrogate models, fine-tuning distance parameters, optimising sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
Figure 3: Key strategies for enhancing LIME, including improvements in sampling, distance parameters, surrogate models, and optimisation.
2.3 Explainable AI (XAI) for Large Language Models
Large Language Models (LLMs) have revolutionised natural language processing, enabling text generation, translation, and question answering tasks. However, their opaque “black-box” nature raises concerns regarding interpretability, fairness, and trustworthiness [72, 73]. Explainability in LLMs is crucial for ensuring accountability, enhancing user trust, and mitigating biases. This literature review explores existing research on explainability methods, applications, and challenges in LLMs.
The conceptual framework for explainability in LLMs includes four key dimensions: Post-hoc Explainability, Mechanistic Interpretability, Human-centric Explainability, and Applications and Challenges. These categories reflect the main approaches and practical considerations in enhancing the transparency and trustworthiness of LLMs.
After training, post-hoc explainability methods aim to interpret LLMs, providing insights into their decision-making processes [72, 74]. Zhao et al. [72] present a taxonomy of explainability techniques for LLMs, categorising approaches into local and global explanations, and discussing limitations and future research directions. Similarly, Bender et al. [73] highlight the risks of opaque LLMs, including biases and misinformation, and call for increased transparency and accountability. Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [75, 76]. Amara et al. [77] recently introduced ConceptX, a concept-level attribution method highlighting semantically meaningful input phrases through a coalition-based Shapley framework, offering robust and human-aligned explanations.
Mechanistic interpretability seeks to uncover how LLMs process information internally [78, 79, 80]. Olah et al. [78] propose reverse-engineering techniques to dissect internal representations. Nanda et al. [79] examine how LLMs develop abstract representations and generalisation, introducing progress measures for interpretability research. Recent work by Anthropic [80] demonstrates the role of mechanistic interpretability in understanding and controlling model behaviour. Additionally, the SEER framework [81] introduces self-explainability mechanisms to enhance the interpretability of internal representations in LLMs. The CELL framework [82] further advances mechanistic interpretability by integrating concept-based explanations directly into the training and representation learning of language models.
Recent advancements from the Transformer Circuits framework have further enriched mechanistic interpretability. Using causal scrubbing, Meng et al. [83] introduced Attribution Graphs to uncover emergent structures within neural networks. In contrast, Nanda et al. [84] investigated the scaling of monosemanticity, extracting interpretable features from state-of-the-art models. Additional recent contributions explore specific tasks such as relevance estimation [85], emotion inference [86], and geospatial reasoning [87]. These works leverage techniques like activation patching and sparse autoencoders to reveal LLM internals. Sparse interpretability methods are further reviewed comprehensively in [88]. Neuroscientific perspectives have also been applied, using dynamical systems theory to model token-level trajectories within transformers [89].
Human-centric explainability focuses on making LLM outputs understandable to non-experts [90, 91, 92]. Ji et al. [90] investigate hallucination in LLM outputs and propose techniques for detecting and mitigating incorrect responses. Krause et al. [91] assess reasoning abilities compared to human performance, and Martens et al. [92] explore counterfactual explanations to improve human understanding. Recent works extend this perspective by enabling interactive explanations of black-box models through fine-tuned LLMs [93] and by benchmarking explanation quality in clinical domains [94]. These approaches improve the accessibility and transparency of LLM outputs in real-world decision-making.
Explainability enhances debugging, bias detection, regulatory compliance, and user trust in LLMs [95]. Challenges remain, including defining “meaning” in LLM-generated outputs [95], managing the ethical and societal implications of black-box models, and balancing performance with interpretability. Recent work has demonstrated SMILE’s applicability beyond language, such as in instruction-based image editing [96], highlighting its modality-agnostic nature. New efforts address rare concept discovery via tailored sparse autoencoders [97] and propose hybrid architectures like concept bottleneck models [98] and symbolic program compositions [99] to enable more transparent reasoning. Future research should emphasise standardised benchmarks, hybrid mechanistic and human-centric methods, and integrating explainability into AI governance frameworks.
3 Problem Definition
This section introduces the fundamental concepts and mathematical framework necessary to understand our approach. We begin by defining key notation and problem settings, followed by a discussion of measuring the impact of perturbations, the surrogate model used for approximation, and a theoretical justification for its fidelity [13, 100, 101]. Finally, we outline a structured workflow for implementing a Wasserstein-based LIME surrogate, ensuring interpretability in the context of large language models [102].
3.1 Notation and Problem Setting
Setting the Stage.
In this study, we aim to develop a local surrogate model that explains the behaviour of a black-box model $\pi^{(n)}$ when provided with input text prompts. The basic notations are as follows:
  • Input: $x$, representing the original prompt.
  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, small modifications of $x$ used to explore the neighbourhood around the original input.
  • Model outputs: $\pi^{(n)}(y\mid x)$, the model’s probability distribution over possible outputs $y$ for the given input $x$.
3.2 Input-Level Distance
To measure how different each perturbed input $\hat{x}_{j}$ is from the original prompt $x$, we compute the semantic distance using Word Mover’s Distance (WMD) [103]:
$\delta_{x_{j}}:=\mathrm{WMD}(x,\hat{x}_{j}).$ (1)
Equation 1 quantifies the semantic dissimilarity between $x$ and $\hat{x}_{j}$, ensuring that closer perturbations are treated as more relevant.
3.3 Output-Level Distribution Shift
To assess how the model’s outputs change when the input is perturbed, we compute the Wasserstein distance between the output distributions for $x$ and $\hat{x}_{j}$ [101]:
$\Delta(x,\hat{x}_{j}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x}_{j})\right).$ (2)
Equation 2 captures the magnitude of the change in the model’s behaviour due to input perturbations.
3.4 Weighting via Input Similarity
Since not all perturbations are equally meaningful, we prioritise those that are semantically closer to the original input. We define a Gaussian-based weighting scheme [13]:
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right).$ (3)
As shown in Equation 3, perturbations with smaller semantic distances $\delta_{x_{j}}$ receive higher weights, reflecting their greater relevance.
3.5 Fitting a Local Surrogate
Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. The surrogate model $h_{\theta}$ adopts a linear form:
$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j}.$ (4)
The parameters $\theta$ are learned by minimising the weighted squared error between the surrogate’s predictions and the observed distribution shifts:
$\min_{\theta}\sum_{j}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2}.$ (5)
Equation 5 ensures that the surrogate closely approximates the black-box model’s behaviour within the local neighbourhood around $x$.
3.6 Theoretical Justification (Lipschitz Smoothness)
Shift Function Definition.
We formally define the shift function:
$f(x,\hat{x}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x})\right).$ (6)
Assuming that $\pi^{(n)}$ is Lipschitz continuous with respect to its input, the shift function $f$ is also Lipschitz [104]:
$|f(x,\hat{x})-f(x^{\prime},\hat{x}^{\prime})|\leq L\cdot\left(\|x-x^{\prime}\|+\|\hat{x}-\hat{x}^{\prime}\|\right),$ (7)
where $L>0$ is a Lipschitz constant (Equation 7). This smoothness justifies the use of a local linear surrogate model to approximate $f$.
Relation to Total Variation.
The Kantorovich-Rubinstein duality [105] relates the Wasserstein distance and Total Variation (TV) distance:
$W(p,q)\leq\text{diam}(\mathcal{Y})\cdot\mathrm{TV}(p,q),$ (8)
where:
$\mathrm{TV}(p,q)=\frac{1}{2}\sum_{y\in\mathcal{Y}}|p(y)-q(y)|.$ (9)
Equations 8 and 9 indicate that small changes in the input resulting in small TV distances also imply bounded Wasserstein shifts.
3.7 Linear Surrogates Approximate Lipschitz Functions
Due to the Lipschitz property, any smooth function $f$ can be locally approximated by a linear model in a small neighbourhood $U$:
$f(x^{\prime})\approx f(x)+\nabla f(x)\cdot(x^{\prime}-x).$ (10)
This principle, expressed in Equation 10, underlies the surrogate modelling approach adopted by LIME and related methods [13].
3.8 Putting It All Together
The complete Wasserstein-based LIME surrogate process consists of the following steps:
  1. 1.
    Generate perturbations $\{\hat{x}_{j}\}$ near $x$.
  2. 2.
    Compute semantic distances $\delta_{x_{j}}$ (Equation 1) and output-level shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
  3. 3.
    Assign weights $w_{j}$ using Equation 3.
  4. 4.
    Extract feature vectors $z_{j}$.
  5. 5.
    Fit a local surrogate model $h_{\theta}$ by minimising the loss (Equation 5).
  6. 6.
    Use $h_{\theta}$ to interpret how input perturbations affect the model’s output distribution.
This framework provides a rigorous, interpretable method for understanding how small changes to an input affect the outputs of complex models, particularly in applications involving large language models [102, 13, 101, 105].
4 PROPOSED METHOD
The proposed approach enhances the interpretability of large language models for text generation by analysing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behaviour [8, 102].
In the context of classification tasks, Fig. 4 illustrates SMILE, a tool designed to explain model predictions by isolating critical features that drive decision-making [12]. SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or logical sections for text, and creates perturbations by selectively modifying these components. These perturbations reveal how different input elements contribute to the final prediction.
Figure 4: SMILE flowchart for explaining image classification [12]
SMILE records the model’s predictions for each perturbed input and calculates similarity scores between the original and modified inputs using metrics such as the Wasserstein distance [100, 101]. These scores are transformed into weights using a kernel function, enabling a balanced weighting scheme. The results are used to train a weighted linear regression model that identifies the most influential input components.
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig. 5. Instead of perturbing the generated image, as in SMILE, we modify the instruction by including or excluding certain words. We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between these outputs and the one produced by the original text prompt [103]. This distance is a similarity measure, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
These similarity scores are then used as the outcome variable in a weighted linear regression model, with weights derived from the text distances and perturbations [106]. This regression model helps determine how each word affects the generated text. The coefficients derived from this model quantify the influence of each word, which we use to create a visual heatmap highlighting the most influential words in the text prompt. This heatmap provides an intuitive visual representation of the elements of the prompt that contribute the most to the generated output.
Using this enhanced three-step process, which builds on SMILE’s methodology [12], we aim to help users understand how particular elements of text prompts impact text generation. This technique improves user control and predictability in text generation models by providing insightful information about which elements of textual input, such as stylistic instructions or descriptive terms, most impact the final output.
The first step employs a text generation model based on input text. The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output. The text generation model is expected to produce responses similar to the original response that vary only based on critical components of the input texts [107].
Different text permutations are utilised from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into individual words. The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [108].
While the text generation model faces challenges in tasks such as counting objects and spatial reasoning, these issues can be mitigated by carefully selecting perturbation texts and restricting the input domain for perturbations [109]. The text generation models enable the proposed method to create responses that differ from the original text solely in terms of keywords related to the perturbation texts.
We use the Wasserstein distance to quantify these differences, comparing the distributions in both text embeddings [110, 100]. This metric is employed because it provides a more comprehensive measure of the differences between text embeddings compared to other ECDF-based distances, such as Kolmogorov–Smirnov or Cramér–von Mises. Although these methods focus on specific aspects, the Wasserstein distance captures both broad and subtle changes in the data [110, 101].
Finally, by applying linear regression, the space of input prompts is linked to the output space of the text generation model. As a result, the model’s behaviour can be interpreted by examining the regression coefficients [111, 106].
4.1 Text Generating and perturbed prompts
To begin, we generate variations of the original text prompt, known as “perturbed prompts”, to examine how these subtle changes affect the output of a text generation model [112, 113]. The original text is first broken down into individual words. Multiple text versions are then created by selectively including or excluding specific words. Each perturbed prompt is then provided to the text generation model, which produces a unique output for each variation. In this process, the output generated from the original prompt is the reference, or baseline output, against which all other outputs are compared.
For the original text prompt $p_{\text{org}}$, the text generation function $\phi_{\text{gen}}$ generates a corresponding output text $t_{\text{org}}^{\prime}$, as shown in Eq. 11:
$t_{\text{org}}^{\prime}=\phi_{\text{gen}}(p_{\text{org}})$ (11)
Similarly, for each perturbed prompt $p_{\text{pert}}^{i}$ within the set $S_{\text{pert}}$, the text generation model produces a corresponding output text $t_{\text{pert}}^{\prime i}$, according to Eq. 12:
$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in S_{\text{pert}}.$ (12)
Here, $p_{\text{org}}$ represents the original prompt, $p_{\text{pert}}^{i}$ describes a perturbation generated from the original text, and $\phi_{\text{gen}}$ is the text generation model function used to produce outputs $t^{\prime}$ that reveal the influence of prompt variations. The set of all perturbed prompts is denoted by $S_{\text{pert}}$ [13, 108].
4.2 Creating the interpretable space
In order to map the output space of the text generation model to a one-dimensional space, the Wasserstein distance between each text generated from the perturbed prompts and the original prompt (baseline output) is computed as a measure of similarity [100, 101]. In all cases, the input prompt remains consistent. Word Mover’s Distance (WMD) is employed to extract meaningful embeddings from the texts, allowing the focus to remain on the primary semantic components of the generated text [103]. This reduces noise in the distance estimation, enabling the accurately calculated distance to serve as the output for the linear regression model.
The distance between the outputs generated by the perturbation texts and the original prompt is given by Eq. 13:
$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{p}\right)^{\frac{1}{p}},\quad\forall i\in S_{\text{pert}}.$ (13)
In the above relation, the $\delta_{\text{WMD}}$ function represents the embedding extraction model applied to the texts before calculating the distance [103]. In Eq. 13, $n$ represents the number of features, $p$ denotes the norm order, and $W(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ demonstrates the Wasserstein distance between the outputs produced by the original and perturbed prompts.
Using p-values is crucial for validating the significance of the observed Wasserstein distances (WD) between texts generated under different prompt perturbations. By calculating a p-value, we test whether the observed WD could have occurred by chance, determining if the perturbation-induced changes are meaningful [114]. To compute this p-value, we use the Bootstrap algorithm [115, 116], which estimates the probability of observing a WD as large as the observed one. This involves generating distributions of WDs through repeated sampling and comparing each sampled WD with the observed value. If a small proportion of sampled WDs exceed the observed WD, the resulting p-value confirms that the differences are statistically significant, supporting the validity of the perturbation effects.
The Wasserstein-$p$ metric allows us to adjust WD’s sensitivity to different text variations by fine-tuning the norm order $p$. Smaller $p$ values emphasise local differences, while larger $p$ values capture broader distributional shifts [101]. By testing various $p$-values, we identify the norm that best captures meaningful variations caused by prompt changes, enhancing the reliability and interpretability of the results [101].
When analysing Wasserstein distances, measures with high p-values, indicating non-significant differences, are excluded from the procedure [114]. As described in Algorithm 1, the Bootstrap algorithm calculates these p-values for all relevant ECDF-based distance measures, including WD. The algorithm computes the WD and its associated p-value.
For a univariate example, let X and Y be the inputs. The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y. In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed. If boostWD exceeds the observed WD, a counter (bigger) is incremented. After completing all iterations, the p-value (pVal) is calculated as the proportion of iterations where boostWD is greater than the observed WD, as shown in Algorithm 1.
Input: Two sets of text-based distances: $X$ and $Y$
Output: Observed Wasserstein Distance (WD) and corresponding p-value (pVal)
1
[…]
6 $n\leftarrow\texttt{LX}+\texttt{LY}$;
7 $\texttt{bigger}\leftarrow 0$;
8
9for $i\leftarrow 1$ to MaxItr do
10       Draw a random sample $e$ of size LX from $XY$;
11       Draw a random sample $f$ of size LY from $XY$;
[…]
14             $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
15
16       end if
17
18 end for
[…]
Algorithm 1 Bootstrap Algorithm for Wasserstein Distance (WD) P-Value Calculation
Figure 5: SMILE flowchart for Large Language Models
When generating perturbation texts derived from the original text, the keywords of the original text play a crucial role in maintaining certain levels of similarity and fidelity between the perturbation texts and the source content. This is important when the interpretable model is trained on perturbed texts as input. Therefore, a similarity criterion is used to measure the degree of similarity between each perturbation text and the original, and this value is employed as a sample weight in the interpretable model. Eq. 15 calculates this similarity, and a Gaussian kernel is applied to normalise it into the range $[0,1]$ [103].
$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\,d(w_{k},w_{l})$ (14)
subject to $\sum_{l=1}^{m}T_{kl}=p_{k},\quad\forall k\in S_{\text{org}}$
$\sum_{k=1}^{n}T_{kl}=q_{l},\quad\forall l\in S_{\text{pert}}$
$T_{kl}\geq 0$
$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\exp\left(-\left(\frac{C_{i}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma^{2}}\right)^{2}\right),\quad\forall i\in S_{\text{pert}}$ (15)
The function $\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ quantifies the semantic distance between the perturbation text and the original. The resulting weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is obtained by applying a Gaussian kernel to this distance [106], allowing the surrogate model to assign higher importance to more semantically similar inputs.
4.3 Developing the Interpretable Surrogate Model
We are innovative in mapping the output space of the text generation model to a one-dimensional distance scale, accompanied by corresponding perturbation texts for the calculated distances [117]. This method enables us to train an interpretable model on the mapped output space of the text generation model, effectively demonstrating how input words influence text generation.
In our linear regression model, the vectors of the perturbation texts are critical, as they are treated as independent variables. Following the extraction of text embeddings, the distance between the outputs of these perturbation texts and the corresponding original text output is regarded as the response variable. Additionally [118], the degree of similarity between each perturbation text and the original text is assumed to represent the weight of the samples [102]. Eq. 16 outlines the square loss objective function for interpretable weighted linear regression:
$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=$ (16)
$\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\times\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$
Here, $\phi_{\text{gen}}$ represents the text generation model, and $\delta_{\text{WMD}}$ is the Word Mover’s Distance function applied to compute similarity between generated outputs. The weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is derived using a Gaussian kernel function and reflects the degree of similarity between the original prompt and each perturbed prompt. The term $\hat{f}_{t}(p_{\text{pert}}^{i})$ is a predicted value from the linear regression model.
After feature extraction, the $\hat{f}_{i}(p_{\text{pert}}^{i,01})$ function is a linear regression model trained over the one-hot vector space of perturbation texts $(p_{\text{pert}}^{i,01})$ and the calculated distances [102].
4.4 Evaluation Metrics
In our investigation, we adopt a suite of evaluation metrics inspired by foundational work provided by Google [119], emphasising the multifaceted nature of assessing explainable models. This work highlights the significance of metrics such as accuracy, stability, fidelity and consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models. The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [119].
Figure 6:
4.4.1 Accuracy
Accuracy measures how well the model’s output aligns with the expected results (ground truth). Specifically, we compare the model’s attribution scores (attention to text elements) against ground truth labels that identify the most relevant text elements for specific generated text segments [120].
To quantify this, we use the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) [121]. AUC reflects the model’s ability to rank relevant elements (from the ground truth) higher than irrelevant ones:
  • AUC $\sim$1: The model effectively distinguishes relevant elements, closely matching human-identified ground truth.
  • AUC $\sim$0.5: The model’s ranking is random, showing poor alignment with the ground truth.
For example, Fig. 6 shows that the ground truth identifies the words “meaning” and “life” as the most relevant text elements for generating the corresponding text output. The metric AttAUC evaluates the model’s alignment with the ground truth, where:
  • AttAUC = 1.0: Perfect alignment.
  • Lower AttAUC (0.8): Less accurate [122].
The heatmaps produced by the model visually represent these attention scores, as shown in Fig. 6, where darker red shades indicate higher attribution to specific text elements. An accurate model should correctly assign higher attention to the text elements that most influence the generated text output, improving the interpretability and reliability of the model [120].
4.4.2 Stability
Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [123]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. Stability is a key property for fostering trust in explainable artificial intelligence (XAI) systems [14]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [15].
In the example text, the attention scores are visualised for two slightly different text prompts. The heatmaps show how the model assigns attention to each word. A stable model, as demonstrated in Fig. 6, assigns similar importance to the words make and rainy in both prompts, even though the token  was added in one case. This consistency is critical for ensuring the model behaves predictably across minor input variations [16]. Without such stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [13].
We use the Jaccard index to quantify stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 17:
$\text{Jaccard}(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (17)
This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater stability and similarity in explanations [123]. A higher Jaccard index confirms the model’s ability to yield stable outputs, thus providing consistent and reliable interpretations [6]. By ensuring stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [14].
4.4.3 Consistency
In evaluating the consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times. Consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [13].
As shown in Fig. 6, when the input phrase “what is the meaning of life?” was fed into the model multiple times, the output remained stable across iterations. This consistency is crucial for applications where repeatable and reliable results are required. In our case, the model consistently assigns higher weights to the words “meaning” and “life”, demonstrating that it understands the importance of these words in generating the desired output [6]. Such behaviour underscores the model’s ability to focus on the semantic core of the input, a trait essential for maintaining coherence in practical use cases [15].
In practical applications—such as automated content generation, interactive systems, or scenario simulations—users depend on models to produce results that do not fluctuate unexpectedly. Any inconsistencies could undermine user trust and compromise the system’s overall effectiveness. Our model enhances user confidence by demonstrating consistent outputs and establishes a foundation for further optimisation and integration into more complex systems. This ability to maintain consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [14].
4.4.4 Fidelity
Fidelity is a metric to evaluate how well an explainable model aligns with a black-box model. The predictions of the black-box model, denoted as $f(x_{i})$, and the explainable model, $g(x_{i})$, for the $i$-th data perturbation are compared. Various loss functions and coefficient metrics quantify the similarity between the models’ predictions, measuring how effectively the explainable model approximates the black-box model.
The coefficient of determination $R^{2}$ is an accuracy-based metric that measures how well the explainable model correlates with the black-box model [124], and is defined in Eq. 18:
$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}(X_{i}))^{2}}$ (18)
where $\bar{f}(X_{i})$ represents the average of $f(X_{i})$. Values of $R^{2}$ close to 1 indicate a strong alignment between the two models.
For scenarios where weighting is necessary, the weighted coefficient of determination $R_{w}^{2}$ provides an alternative version [16, 125], defined as Eq. 19:
$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f_{w}}(X_{i}))^{2}}$ (19)
where $\bar{f_{w}}(X_{i})$ represents the weighted average of $f(X_{i})$. These accuracy-based metrics, $R^{2}$ and $R_{w}^{2}$, offer insight into how well the explainable model approximates the black-box model.
To account for both sample size and the number of features used in the local explanation, the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ [16] is used, as shown in Eq. 20:
$\hat{R}_{w}^{2}=1-(1-R_{w}^{2})\left[\frac{N_{p}-1}{N_{p}-N_{s}-1}\right]$ (20)
To measure the error or risk between the models’ predictions, error-based metrics, including the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE), are used [126]. These metrics quantify the weighted differences between the predictions of the explainable model and those of the black-box model. Lower values indicate better alignment. WMSE and WMAE are defined in Eqs. 21 and 22:
$\text{WMSE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}w_{i}}$ (21)
$\text{WMAE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot|f(X_{i})-g(X_{i})|}{\sum_{i=1}^{N_{p}}w_{i}}$ (22)
Here, $w_{i}$ denotes the weight for each data point $i$, $f(X_{i})$ represents the predictions from the black-box model, $g(X_{i})$ represents the predictions from the explainable model, and $N_{p}$ is the total number of data points.
Additionally, the mean $L_{1}$ and mean $L_{2}$ losses, defined below, are commonly used to assess prediction discrepancies [127, 16], as shown in Eqs. 23 and 24:
$L_{1}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (23)
$L_{2}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (24)
These error-based metrics, including MSE, MAE, $L_{1}$, and $L_{2}$, quantify the differences between the black-box and explainable models’ predicted scores, reflecting the risk of prediction discrepancies.
We compute fidelity across various scenarios. Fidelity is measured by comparing the predictions of the explainable model to those of the black-box text generation model. We analyse how different perturbations in the input text affect fidelity scores and how various distance metrics for text-to-text comparisons perform when using linear regression and Bayesian ridge as surrogate models.
5 Experimental Results
This section evaluates the proposed method’s ability to enhance explainability in instruction-based text generation. We analyse how the structure of textual prompts affects interpretability, how variations in input text influence the clarity of explanations, and how the proposed solution performs across diverse scenarios.
Experiments were conducted using a range of datasets and four state-of-the-art text generation models, OpenAI GPT [1], LLaMA [5], and Claude-AI [128] to ensure a comprehensive assessment.
5.1 Qualitative Results
The qualitative evaluation demonstrates the proposed framework’s interpretability using scenario-based testing and visualisations. We designed diverse testing scenarios by varying textual prompts, including changes in sentence complexity, rephrasing, and the inclusion of domain-specific terminology. These scenarios examine how the model interprets and generates outputs based on nuanced variations in textual inputs.
5.1.1 Input Prompt Explainability
The proposed framework applies scenario-based testing across multiple prompt variations to evaluate the interpretability of the model’s behaviour regarding different inputs. These variations include rephrased questions, descriptive modifiers, and changes in word choice. The system generates a text heatmap highlighting individual words’ contribution to the model’s output for each input, with colours indicating the importance scores.
Figure 7 shows how varying the input prompt’s phrasing or focus leads to different word importance patterns, demonstrating the model’s sensitivity to linguistic changes.
Figure 7: Text heatmaps showing word-level importance for various input prompts. The colour intensity represents the influence of each word on the model’s output.
5.1.2 Reasoning Path and Token Attribution
Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response. Figure 8 presents an example where the SMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram. This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models. As shown in Figure 8, SMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [129].
Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the SMILE framework. The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates SMILE’s token-level importance heatmap for the same prompt. [129]
5.1.3 Word-Level Attribution Analysis on MMLU Using SMILE
To evaluate the interpretability of ChatGPT using the SMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [130]. This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities. For this study, we accessed a pre-formatted version of the dataset available on Kaggle [131] to facilitate prompt construction and experimentation.
Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision. Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter. These visualisations demonstrate SMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
Figure 9: Attribution heatmaps generated by SMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
5.2 Quantitative Results
In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to asses our explainability method across different text generation models and different scenarios.
5.2.1 Accuracy
To evaluate the performance of instruction-based text generation models, we tested ten different prompts and their variations through 64 perturbations. Heatmaps were generated to display the weights of each keyword in the prompts across various models, including OpenAI GPT, LLAMA, and Claude-AI. A ground truth was defined, assigning weights of 1 to critical keywords in the prompts and 0 to less relevant or auxiliary words. These ground truth values were compared with the extracted heatmaps to assess the interpretability and accuracy of each model.
To quantify accuracy, we employed multiple metrics, including Attention Accuracy (ATT ACC), F1-Score for Attention (ATT F1), and Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC). The results were averaged across the ten different prompts, and each model’s performance is summarised in Table 1, showcasing their effectiveness based on these metrics.
Table 1: Performance metrics (ATT ACC, ATT F1, and ATT AUROC) for various
models, evaluated under two conditions: different prompts and different images.
Model name ATT ACC ATT F1 ATT AUROC
OpenAI-GPT 0.70 0.59 0.84
LLaMA 0.76 0.40 0.84
Claude-AI 0.82 0.67 0.88
5.2.2 Stability
To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap. For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word. To evaluate stability, we added the token ‘‘ at the end of each prompt as input text, extracted the coefficients again, and computed the Jaccard index to compare the sets of coefficients before and after the perturbation. Finally, we calculated the average Jaccard index for each model, and the results are presented in Table 2.
Table 2: Stability across different models for ten prompts and 30 perturbations
Model name Jaccard Index
OpenAI-GPT 0.62
LLaMA 0.45
Claude-AI 0.44
5.2.3 Consistency
To compute consistency, we ran the same code with 64 perturbations for each prompt over 10 iterations, ensuring that the outputs remained consistent across these repetitions for text generation models such as OpenAI GPT, LLAMA, and Claude-AI, as shown in Table 3. We computed variance and standard deviation metrics for each word coefficient. This step is essential to confirm that the model’s predictions are not random or heavily dependent on initialisation factors but are instead deterministic and robust.
Table 3: Consistency metrics for different models on the prompt
“What is the meaning of life?”
Model name Variance Standard Deviation
OpenAI-GPT 0.0000 0.0046
LLaMA 0.0048 0.0663
Claude-AI 0.0011 0.0331
5.2.4 Fidelity for Different Instruct image editing diffusion based Models (IED)
[…]
Table 4: Fidelity metrics for different models on the prompt
“What is the meaning of life?”
[…]
OpenAI-GPT 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
LLaMA 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
Claude-AI 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
5.2.5 Fidelity Across Different number of Text Perturbations
Tables 56, and 7 present the results of fidelity computations for varying numbers of perturbations across the OpenAI GPT, LLAMA, and Claude-AI models, respectively. These results are derived using a weighted linear regression as the surrogate model and the Wasserstein distance for computing distances. The evaluation metrics include weighted mean squared error (WMSE), weighted mean absolute error (WMAE), coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$), and average loss metrics such as mean-L1 and mean-L2.
Table 5: Performance metrics for different numbers of perturbations in OpenAI GPT
[…]
32 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
64 0.0617 0.3406 0.1695 0.1756 0.0526 0.2712
128 0.0329 0.4468 0.1385 0.1519 0.0378 0.4193
256 0.0116 0.4276 0.0675 0.0872 0.0163 0.4138
Table 6: Performance metrics for different numbers of perturbations in LLAMA
[…]
32 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
64 0.0240 0.4844 0.1284 0.1433 0.0382 0.4301
128 0.0085 0.2430 0.0638 0.0612 0.0096 0.2055
256 0.0058 0.2387 0.0605 0.0667 0.0118 0.2203
Table 7: Performance metrics for different numbers of perturbations in Claude-AI
[…]
32 0.0368 0.7209 0.1495 0.2262 0.0733 0.6539
64 0.0115 0.3708 0.0564 0.1036 0.0370 0.3046
128 0.0089 0.1506 0.0535 0.0920 0.0192 0.1085
256 0.0191 0.2670 0.0902 0.0873 0.0163 0.2494
5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
[…]
Table 8: Fidelity results for different distance measures with 30 perturbations
WLR Fidelity Metrics
[…]
Cosine Cosine 0.0172 0.3151 0.0659 0.1277 0.0412 0.1508
Cosine WD 0.0216 0.4197 0.0899 0.1332 0.0385 0.2805
WD WD 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
WD Cosine 0.0048 0.4026 0.0329 0.0871 0.0296 0.2593
WD+C WD+C 0.0349 0.5349 0.1589 0.3050 0.1468 0.4233
BayLIME
Cosine Cosine 0.0200 0.2285 0.0499 0.1067 0.0464 0.0434
Cosine WD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
WD WD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
WD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
WD+C WD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
5.2.7 Computation complexity
We apply our model to three instruction-based text generation frameworks: OpenAI GPT, LLaMA, and Claude-AI. We also evaluate the execution time of different explainability methods: LIME, SMILE, and Bay-LIME. Each method is tested with 60 perturbations on the same hardware to ensure consistency in computational conditions.
The models exhibit notable differences in execution time, primarily due to how each explainability method computes similarity and assigns weights. SMILE, in particular, employs the Wasserstein distance to quantify text dissimilarities, making it substantially more computationally expensive than both LIME and Bay-LIME. As reported in Table 9, SMILE consistently incurs higher execution times across all three model frameworks. This pattern also holds across other frameworks, indicating that SMILE’s computational overhead is a systemic characteristic of its underlying methodology.
Wasserstein distance is known for capturing subtle differences between text embeddings, making it especially effective in high-dimensional semantic spaces [101]. However, this precision comes at a cost. Computing the Wasserstein-2 distance involves operations on covariance matrices with time complexity proportional to $O(d^{3})$, where $d$ is the embedding dimensionality. When applied across $N$ samples, the total complexity becomes $O(Nd^{3})$. In contrast, LIME relies on cosine similarity, which has a much lower computational cost of $O(Nd)$.
This disparity arises because Wasserstein-2 distance requires solving matrix optimisation problems over $d$-dimensional Gaussian distributions [132]. While this added complexity leads to more stable and expressive explanations, as demonstrated by SMILE, it significantly increases runtime. Conversely, LIME avoids this overhead by utilising simpler vector-based similarity measures.
In summary, SMILE offers greater fidelity and stability but at a higher computational cost. Its use of the Wasserstein distance makes it ideal for tasks where precision and interpretability are essential, even if that entails longer execution times. LIME and Bay-LIME, while faster, may trade off some robustness in favour of efficiency.
Table 9: Execution times (in seconds) for each framework and explainability method with 60 perturbations
Framework LIME SMILE Bay-LIME
OpenAI-GPT 156.82 170.70 161.43
LLaMA 113.00 118.99 116.40
Claude-AI 355.09 372.95 349.88
5.2.8 Optimised Instructions from Google Research
To further explore how instruction phrasing influences interpretability, we applied SMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [133]. These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
We first applied SMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem. ” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, SMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using SMILE with GPT-3.5.
Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using SMILE with GPT-4.
5.2.9 Token-Level Attention for Different Sentence Structures
To investigate how sentence structure affects token-level attention in large language models (LLMs), we analyzed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts. These prompts ranged from questions (e.g., “What is the meaning of life?”), commands (e.g., “Please give me the meaning of life”), to declarative formulations (e.g., “You must explain the meaning of life”).
For each sentence, we extracted the attention values directly from the final layer of the LLM. We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualized using a box plot (see Figure 12), where each point represents a unique sentence. Full-sentence prompts are viewable through hover text, allowing qualitative comparison of sentence structure with attention distribution.
The analysis reveals noticeable variability in how these tokens are weighted depending on syntactic form. For instance, in imperative constructions, attention to “life” tends to dominate, whereas in interrogative or abstract prompts, “meaning” often receives comparable or higher attention. This suggests that sentence formulation can subtly steer the internal focus of the model, even when the semantic content remains consistent.
These findings underscore the sensitivity of LLMs to surface-level structure in shaping token-level interpretability, offering insight into both model behaviour and prompt engineering strategies.
Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures. Hovering over each point reveals the full sentence associated with that score.
6 Limitations
While SMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, even when the semantic meaning remains largely intact. This sensitivity challenges the assumption of local smoothness required for regression-based attribution and can reduce explanation stability.
2) Some LLMs exhibit hallucination, producing fluent but factually incorrect or irrelevant outputs, especially when presented with ambiguous or syntactically degraded perturbations. This introduces noise into the attribution signal and affects explanation reliability.
3) The computational cost of generating and processing many perturbations (e.g., 1024 per input) is significant. Commercial APIs can lead to high inference costs and latency. For local models, substantial GPU resources are required, limiting practical scalability.
4) Some models, particularly research-stage or open-source LLMs, do not provide API access and require manual local deployment.
7 Conclusion
Interpretability remains a cornerstone of responsible AI, particularly in the context of large language models (LLMs) and instruction-based text generation. As AI systems become increasingly embedded in domains such as business, education, healthcare, and public services, it becomes essential for users to understand how these models generate outputs from their prompts. Without such insight, trust in AI systems diminishes, limiting adoption in real-world workflows. This underscores the need for transparent and interpretable solutions to demystify the behaviour of these black-box models.
Interpretability is not only a research priority but is likely to become a legal and operational requirement in alignment with emerging AI regulations and ethical standards. Organisations that fail to address interpretability may encounter regulatory hurdles and reduced user acceptance.
This work introduced a novel interpretability framework tailored for instruction-based text generation models. Leveraging SMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output. Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
Whereas most existing interpretability methods highlight important tokens in the model’s output [13, 134], SMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that SMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma. The resulting heatmaps and attribution scores intuitively understand model behaviour, empowering users to control better and refine prompt design.
This work lays the groundwork for more transparent and trustworthy natural language systems. Future directions include extending SMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
8 Future Works
  1. 1.
    Extension to Mathematical and Equation-Based Text Generation: Future work will explore the application of the proposed interpretability framework to models focused on symbolic reasoning and math-related tasks, such as equation solving, theorem generation, and formal derivation. Instruction-based mathematical models like Minerva and MathGLM require high precision in prompt design, and understanding how prompt components affect symbolic output will be essential in educational and scientific settings.
  2. 2.
    Application to Paraphrase Generation Models: The methodology can be extended to paraphrasing models, where semantic preservation and variation are central. Analysing how individual words influence paraphrased outputs’ style, tone, or fidelity can support better user control in writing assistance, content rephrasing, and translation tools.
  3. 3.
    Adaptation to Multi-Turn and Context-Aware Generation: Another direction is to adapt the framework for multi-turn and context-aware systems, including dialogue agents. This would involve tracking how prompt changes impact responses’ consistency, coherence, and relevance across conversational sequences.
  4. 4.
    Cross-Task Generalisation: While this work focuses on instruction-based text generation, the interpretability strategy can be generalised to broader NLP tasks, including summarisation, classification, and question answering. Evaluating how different model families respond to prompt perturbations can uncover model-specific behaviours and improve transparency.
  5. 5.
    Benchmarking Interpretability and Prompt Sensitivity: We plan to contribute to standardised evaluation protocols for prompt sensitivity, attribution robustness, and local fidelity. These benchmarks will support consistent assessment across models and help identify best practices for interpretable generation.
Data Availability
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the SMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.