### v2

# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

Abstract
Large Language Models (LLMs) such as GPT, LLaMA, and Claude have achieved remarkable performance in text generation, yet their decision-making processes remain opaque. This lack of transparency poses challenges in domains where trust and accountability are critical. In this work, we introduce gSMILE (Generative SMILE), a generalized and LLM-oriented extension of the previously proposed SMILE framework, for interpreting how these models respond to specific parts of an input prompt.
gSMILE is model-agnostic and operates by introducing controlled perturbations to the input, observing the resulting output changes, and identifying which input words have the most significant impact on the output. This process enables the generation of intuitive heatmaps that visually highlight influential words in the prompt. We evaluate gSMILE across several leading large language models (LLMs) using metrics such as accuracy, consistency, stability, and fidelity, demonstrating its ability to provide precise, robust, and reliable explanations. By making LLMs more interpretable, gSMILE takes an important step toward building more transparent and trustworthy AI systems.
Figure 1: Highlighting the contribution of each word in the input prompt of a Commercial LLM
1 Introduction
Large Language Models (LLMs), such as GPT, Claude, and LLaMA, have significantly advanced the capabilities of natural language processing systems [1, 2, 3]. These models, based on the Transformer architecture [4], are trained on extensive corpora and exhibit remarkable performance across a wide range of tasks, including text generation, summarisation, code completion, and dialogue reasoning [5]. However, despite their widespread adoption, LLMs remain largely opaque in how they make predictions, leading to concerns about transparency, trust, and safety, especially in high-stakes domains such as healthcare and law [6, 7].
The black-box nature of LLMs limits our ability to understand which parts of an input prompt have the most significant influence on the model’s output. This lack of interpretability hinders debugging, ethical assessment, and responsible deployment [8, 9]. Interpretability methods, especially those that operate post hoc and in a model-agnostic fashion, offer a way to open these black boxes by explaining the model’s behavior without requiring access to internal parameters [10].
In this work, we propose gSMILE, a generalized and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [11]. gSMILE interprets model outputs by perturbing input prompts, measuring output shifts using Wasserstein distance, and fitting a local linear surrogate model. The result is a weighted heatmap that visualizes how individual words in a prompt contribute to the model’s response.
We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including fidelity, consistency, stability, and accuracy. Our findings demonstrate that gSMILE yields stable, interpretable, and reliable attributions, rendering it suitable for practical applications in real-world scenarios involving instruction-based generation.
The key contributions of this paper are:
  • Token-level interpretability for LLMs: We present gSMILE, a model-agnostic method that quantifies the influence of each input token on the output of large language models using perturbation-based analysis and Wasserstein distance.
  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behavior.
  • Robust cross-model evaluation: We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, accuracy, fidelity, stability, and consistency that demonstrate its reliability and adaptability.
2 Literature review
This section provides an overview of recent advancements in large language models (LLMs) and their increasing impact across various applications. We also introduce Explainable Artificial Intelligence (XAI), which aims to make AI models more transparent, interpretable, and trustworthy. As LLMs grow in complexity and usage, the need to understand their decision-making processes becomes more critical. Therefore, we highlight the emerging efforts to develop explainable large language models (LLMs), outlining key challenges and recent techniques that aim to improve their interpretability and foster greater trust in their outputs.
2.1 Large Language Models
Large language models (LLMs) have become increasingly prevalent due to significant advancements in deep learning [12], improved access to large-scale datasets, and the growing availability of powerful computing resources [13]. These models are typically trained on massive corpora of text. They are capable of generating fluent, contextually appropriate responses across a wide variety of language-based tasks, including question answering, summarisation, translation, and content generation. Over time, several influential large language models (LLMs) have helped push the field forward, each introducing novel architectural innovations [4], training strategies [14], or alignment techniques [15] that have contributed to their growing capabilities and broader adoption.
Among the most influential models, GPT (OpenAI) [1] stands out as a leading autoregressive transformer-based architecture capable of generating coherent and contextually rich text. It set a new standard for generative language models and introduced the broader public to powerful applications of AI through tools like ChatGPT. Expanding on this paradigm, Google’s T5 [14] reframed all NLP tasks as text-to-text problems, streamlining the transfer of learning across diverse tasks and contributing to more unified and efficient architectures. Palm (Google) [16], with its Pathways design, extended the multitask learning capability at scale, allowing a single model to generalize effectively across different NLP applications.
In parallel, Anthropic’s Claude [17] introduced a new approach centered on safe and aligned AI behavior. Claude utilizes constitutional AI principles to ensure its outputs are consistent with predefined ethical standards, thereby promoting greater transparency and reducing the risk of harm. More recently, Google’s Gemma 2 [18] has focused on logical reasoning and structured data interpretation, offering a competitive option for research-driven and enterprise-level deployment. Meta’s LLaMA [19] series continues to draw attention as a scalable, open-source alternative, providing the research community with accessible tools that deliver high performance across a range of benchmarks.
Community-driven efforts have also enriched the landscape. EleutherAI’s GPT-NeoX [20] was developed to provide an open-source alternative to proprietary GPT models, offering researchers greater flexibility in experimentation and deployment. AI21 Labs’ Jurassic-2 [21] targets commercial-scale applications with a focus on contextual awareness and enterprise integration. Meanwhile, Cohere’s Command R+ [22] integrates retrieval-augmented generation (RAG), enhancing factual consistency and minimizing hallucinations in the generative output — a key concern for the practical adoption of large language models (LLMs).
Additionally, transformer-based models, such as BERT [23], have revolutionized NLP through bidirectional encoding, providing substantial improvements in tasks like question answering, named entity recognition, and sentiment analysis. Efficiency-focused models, such as Microsoft’s Phi-4, highlight a growing trend toward parameter-efficient design without sacrificing core performance, making large language models (LLMs) more accessible for low-resource environments. Mistral and Qwen 2.5 (by Alibaba) continue this direction by offering fine-tuning-friendly models with strong multilingual and domain adaptation capabilities.
Several recent open-source contenders further enrich the ecosystem. DeepSeek-LLM [24] delivers performance comparable to proprietary models while remaining open and extensible. Mixtral and StableLM focus on task efficiency and targeted performance, ensuring their utility in production environments. DeepSeek-R1 has been noted for its emphasis on reasoning, reflecting a broader shift from general language generation to models optimized for critical thinking, chain-of-thought prompting, and interpretability.
Cutting-edge advancements have introduced even more specialized architectures. Qwen2.5-Max utilizes large-scale Mixture-of-Experts (MoE) models to enhance throughput and scalability. OpenAI’s o3-mini emphasizes low-latency, cost-effective deployment while retaining high-quality output — ideal for mobile and edge computing. Perhaps most notably, DeepSeek-V3 represents a significant milestone by achieving GPT-4o-level performance in an open-source setting, suggesting that open-source models may soon rival or surpass closed models in both capability and adoption.
The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications. These models continue to define the state of the art in machine-generated language, with a growing focus on transparency, efficiency, and alignment with human values.
Category Models Description General-Purpose GPT-4 [25], GPT-4o [26], Claude 3 [27], Gemini 1.5 [28], Command R+ [29], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [32], Falcon 180B [33], DeepSeek-V3 [34], OpenChat [35] Customizable models for research applications Retrieval-Augmented Command R+ [29], GPT-4 Bing [25], Claude 3 RAG [27], Gemini Search [28] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [27], Gemini 1.5 [28], GPT-4o [26], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [25], Claude 3 Opus [27], Gemini 1.5 [28], Mixtral [32], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [27], Gemini [28], GPT-4 [25], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [32], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [26], Gemini 1.5 [28], Claude 3 [27], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [43], XLM-R [44], ByT5 [45] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [46], Galactica [47], Legal-BERT [48], StarCoder2 [49], Code LLaMA [50] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [51], Vicuna [52], ChatGLM [53], Baize [54], OpenChat [35] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [55], TinyLLaMA [56], DistilGPT2 [57], GPT-2 [58], LLaMA 2 7B [59] Designed for fast inference and deployment on edge devices
Figure 2: Extended LLM Categories, Models, and Descriptions
2.2 Explainable AI (XAI)
The increasing complexity of machine learning models has made understanding how they make decisions more critical than ever. Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [60, 61].
Explainability methods can be broadly categorized as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [6].
In contrast, post-hoc methods, such as Local Interpretable Model-agnostic Explanations (LIME) [63] and SHAP (Shapley Additive Explanations) [64], are applied after model training to explain the decisions of complex, often “black-box” models like neural networks and ensemble methods. These approaches provide valuable insights without altering the underlying model structure [65]. Post-hoc explainability methods can be classified into global and local explanations.
Global explanations offer an overarching view of the model’s decision-making patterns across an entire dataset, facilitating the identification of trends and biases [62, 66]. In contrast, local explanations focus on specific predictions, providing a detailed breakdown of how individual inputs influence model outputs. LIME and its numerous enhancements fall under local explainability methods and are designed to provide granular, case-specific interpretations [63].
LIME has become one of the most widely used explainability techniques; however, over time, researchers have introduced various improvements to address its limitations and expand its applicability across different applications. As shown in Figure 3, enhancements to LIME can be grouped into four main categories.
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [67], S-LIME [68], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behavior, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimization strategies to enhance the efficiency and stability of LIME. Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [69], TS-MULE for time series data [70], Graph LIME for graph-structured data [71], Sound LIME for audio data [72], and B-LIME for ECG signal data [73] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimizing sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
Other extensions include DSEG-LIME [74], which tailors LIME for segmentation models by generating perturbations at the object level rather than the pixel level, improving interpretability in image segmentation tasks. SLICE [75] introduces structured sampling and clustering to ensure diverse and representative local explanations, reducing redundancy in LIME samples. Stratified LIME [76] improves fidelity by stratifying the input space and generating more targeted local samples for different subpopulations. US-LIME [77] employs uncertainty sampling to guide instance selection, increasing the reliability and faithfulness of local explanations. SS-LIME [78] uses self-supervised learning to identify important regions in the feature space and provide richer explanations. K-LIME [79] incorporates kernel-based methods into LIME’s regression step to model more complex local decision boundaries. Attention-LIME [80] leverages attention mechanisms to prioritize features during the explanation, aligning more closely with deep learning models. BMB-LIME [81] employs a boosted multi-branch approach to combine multiple local models, thereby enhancing consistency and robustness.
These methods aim to enhance interpretability across various domains, including segmentation, survival analysis, multimodal learning, and attention-based contexts. Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [82], LIME-SUP, LORE, ILIME, Aleph LIME, bLIMEy, NormLIME, DLIME, LIMEtree, MPS-LIME, Explain Explore, survLIME, survLIME-inf, CBR-LIME, GuidedLIME, audioLIME, MeLIME, C-LIME, gLIME, DIME, s-LIME, LIMESegment, UnRAvEL-LIME, Kernel-LIME, Specific-Input LIME, ExpLIMEable, GLIME, GMM-LIME, SEGAL, and Sig-LIME, reflecting the diverse and rapidly evolving landscape of post-hoc explainability [83].
Figure 3: Taxonomy of LIME enhancement strategies: categorized into sampling techniques, surrogate model improvements, distance parameter adjustments, and optimization approaches.
2.3 Explainable AI (XAI) for Large Language Models
Large Language Models (LLMs) have revolutionised natural language processing, enabling text generation, translation, and question answering tasks. However, their opaque “black-box” nature raises concerns regarding interpretability, fairness, and trustworthiness [84, 85]. Explainability in LLMs is crucial for ensuring accountability, enhancing user trust, and mitigating biases. This literature review explores existing research on explainability methods, applications, and challenges in LLMs.
The conceptual framework for explainability in LLMs includes four key dimensions: Post-hoc Explainability, Mechanistic Interpretability, Human-centric Explainability, and Applications and Challenges. These categories reflect the main approaches and practical considerations in enhancing the transparency and trustworthiness of LLMs.
After training, post-hoc explainability methods aim to interpret LLMs, providing insights into their decision-making processes [84, 86]. Zhao et al. [84] present a taxonomy of explainability techniques for LLMs, categorising approaches into local and global explanations, and discussing limitations and future research directions. Similarly, Bender et al. [85] highlight the risks of opaque LLMs, including biases and misinformation, and call for increased transparency and accountability. Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [87, 88]. Amara et al. [89] recently introduced ConceptX, a concept-level attribution method highlighting semantically meaningful input phrases through a coalition-based Shapley framework, offering robust and human-aligned explanations.
Mechanistic interpretability seeks to uncover how LLMs process information internally [90, 91, 92]. Olah et al. [90] propose reverse-engineering techniques to dissect internal representations. Nanda et al. [91] examine how LLMs develop abstract representations and generalisation, introducing progress measures for interpretability research. Recent work by Anthropic [92] demonstrates the role of mechanistic interpretability in understanding and controlling model behaviour. Additionally, the SEER framework [93] introduces self-explainability mechanisms to enhance the interpretability of internal representations in LLMs. The CELL framework [94] further advances mechanistic interpretability by integrating concept-based explanations directly into the training and representation learning of language models.
Recent advancements from the Transformer Circuits framework have further enriched mechanistic interpretability. Using causal scrubbing, Meng et al. [95] introduced Attribution Graphs to uncover emergent structures within neural networks. In contrast, Nanda et al. [96] investigated the scaling of monosemanticity, extracting interpretable features from state-of-the-art models. Additional recent contributions explore specific tasks such as relevance estimation [97], emotion inference [98], and geospatial reasoning [99]. These works leverage techniques like activation patching and sparse autoencoders to reveal LLM internals. Sparse interpretability methods are further reviewed comprehensively in [100]. Neuroscientific perspectives have also been applied, using dynamical systems theory to model token-level trajectories within transformers [101].
Human-centric explainability focuses on making LLM outputs understandable to non-experts [102, 103, 104]. Ji et al. [102] investigate hallucination in LLM outputs and propose techniques for detecting and mitigating incorrect responses. Krause et al. [103] assess reasoning abilities compared to human performance, and Martens et al. [104] explore counterfactual explanations to improve human understanding. Recent works extend this perspective by enabling interactive explanations of black-box models through fine-tuned LLMs [105] and by benchmarking explanation quality in clinical domains [106]. These approaches improve the accessibility and transparency of LLM outputs in real-world decision-making.
Explainability enhances debugging, bias detection, regulatory compliance, and user trust in LLMs [107]. Challenges remain, including defining “meaning” in LLM-generated outputs [107], managing the ethical and societal implications of black-box models, and balancing performance with interpretability. Recent work has demonstrated SMILE’s applicability beyond language, such as in instruction-based image editing [108], highlighting its modality-agnostic nature. New efforts address rare concept discovery via tailored sparse autoencoders [109] and propose hybrid architectures like concept bottleneck models [110] and symbolic program compositions [111] to enable more transparent reasoning. Future research should emphasise standardised benchmarks, hybrid mechanistic and human-centric methods, and integrating explainability into AI governance frameworks.
3 Problem Definition
This section introduces the core concepts and mathematical framework underlying our approach. We begin by defining the key notation and problem setup, followed by a formal description of how we quantify the impact of input perturbations. We then describe the use of a surrogate model to approximate model behavior locally and provide theoretical justification based on the Wasserstein distance [63, 112, 113]. Lastly, we outline a structured methodology for implementing a Wasserstein-based surrogate framework tailored to large language models [114].
3.1 Notation and Problem Setting
Setting the Stage.
We aim to construct a local surrogate model that explains the behavior of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt. The goal is to understand how specific parts of the input affect the model’s output distribution. The following notation is used:
  • Input: $x\in\mathcal{X}$, representing the original input prompt.
  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, a set of $J$ perturbed versions of $x$, generated via minor token-level edits.
  • Model Outputs: $\pi^{(n)}(y\mid\hat{x}_{j})$, the model’s probability distribution over output space $\mathcal{Y}$, conditioned on the perturbed input $\hat{x}_{j}$. For the unperturbed input $x$, we write $\pi^{(n)}(y\mid x)$.
3.2 Input-Level Distance
To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [115]:
$\delta_{x_{j}}=\mathrm{WMD}(x,\hat{x}_{j}),$ (1)
Where $\delta_{x_{j}}\in\mathbb{R}_{\geq 0}$ denotes the semantic distance between the original and perturbed prompts.
3.3 Output-Level Distribution Shift
To assess the effect of input perturbation on the model’s behavior, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs:
$\Delta(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
where $y\in\mathcal{Y}$, and $W(p,q)$ denotes the Wasserstein distance between two probability distributions $p$ and $q$.
3.4 Weighting via Input Similarity
Perturbations that are semantically closer to the original input are weighted more heavily. We use a Gaussian-based weighting scheme [63]:
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right),$ (3)
Where $w_{j}\in(0,1]$ denotes the relevance weight of $\hat{x}_{j}$, and $\sigma>0$ is a kernel width parameter that controls the rate of exponential decay.
3.5 Fitting a Local Surrogate
Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. We fit a linear surrogate model of the form:
$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j},$ (4)
where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimizing the weighted squared error.
$\min_{\theta}\sum_{j=1}^{J}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2},$ (5)
Where the loss ensures local fidelity between the surrogate and the observed output shifts.
3.6 Theoretical Justification (Lipschitz Smoothness)
Shift Function Definition.
We define a shift function that quantifies the model’s response to perturbation:
$f(x_{a},\hat{x}_{a})=W\left(\pi^{(n)}(y\mid x_{a}),\pi^{(n)}(y\mid\hat{x}_{a})\right),$ (6)
If $\pi^{(n)}$ is Lipschitz continuous concerning its input, then the function $f$ is also Lipschitz:
$\left|f(x_{a},\hat{x}_{a})-f(x_{b},\hat{x}_{b})\right|\leq L\times\left(\|x_{a}-x_{b}\|+\|\hat{x}_{a}-\hat{x}_{b}\|\right),\forall a,b\in\{1,\dots,N_{pert}\}$ (7)
The notation $x_{a}$ and $\hat{x}_{a}$ represents a specific input and its corresponding perturbation. The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analyzing the Lipschitz continuity or stability across perturbations. $N_{pert}$ denotes the total number of perturbations.
[…]
Relation to Total Variation.
For two distributions $p$ and $q$ over output space $\mathcal{Y}$, the Kantorovich–Rubinstein duality yields:
$W(p,q)\leq\text{diam}(\mathcal{Y})\cdot\mathrm{TV}(p,q),$ (8)
Where the total variation distance ($TV$) is given by:
$\mathrm{TV}(p,q)=\frac{1}{2}\sum_{y\in\mathcal{Y}}|p(y)-q(y)|.$ (9)
This bound highlights that small changes in distribution (in the TV sense) imply bounded Wasserstein distance.
3.7 Linear Surrogates Approximate Lipschitz Functions
Due to the differentiability of Lipschitz functions, the shift function $f$ can be locally approximated by a first-order Taylor expansion:
$f(x^{\prime})\approx f(x)+\nabla f(x)^{\top}(x^{\prime}-x),$ (10)
where $\nabla f(x)^{\top}(x^{\prime}-x)$ denotes the dot product between the gradient of $f$ and the local displacement in input space.
3.8 Putting It All Together
The complete Wasserstein-based LIME surrogate procedure consists of:
  1. 1.
    Generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$.
  2. 2.
    Computing semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
  3. 3.
    Assigning similarity weights $w_{j}$ via Equation 3.
  4. 4.
    Extracting feature representations $z_{j}$ for each perturbation.
  5. 5.
    Fitting a local surrogate model $h_{\theta}$ by minimizing the loss (Equation 5).
  6. 6.
    Using $h_{\theta}$ to interpret the influence of each input feature on model behavior.
This framework offers a principled and interpretable approach to analyzing large language models through localized perturbation and response analysis.
4 PROPOSED METHOD
The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analyzing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behavior [9, 114].
In the context of classification tasks, Fig. 4 illustrates SMILE, a tool designed to explain model predictions by isolating critical features that drive decision-making [11]. SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or “logical sections” for text, and creates perturbations by selectively modifying these components. These perturbations reveal how different input elements contribute to the final prediction.
Figure 4: SMILE flowchart for explaining image classification [11]
SMILE records the model’s predictions for each perturbed input. It calculates distance scores between the original and modified outputs using metrics such as the Wasserstein distance [112, 113]. These distances are transformed into similarity-based weights using a kernel function, enabling a balanced weighting scheme. The results are used to train a weighted linear regression model that identifies the most influential input components.
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig. 5. Instead of perturbing the generated image, as in SMILE, we perturb the input prompt by including or excluding certain words. We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between each generated output and the original output [115]. This distance quantifies the difference in the production, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
These similarity scores are then used as the outcome variable in a weighted linear regression model, with weights derived from the text distances and perturbations [116]. This regression model helps determine how each word affects the generated text. The coefficients derived from this model quantify the influence of each word, which we use to create a visual heatmap highlighting the most influential words in the text prompt. This heatmap provides an intuitive visual representation of the elements of the prompt that contribute the most to the generated output.
Using this enhanced three-step process, which builds on SMILE’s methodology [11], we aim to help users understand how particular elements of text prompts impact text generation. This technique enhances user control and predictability in text generation models by providing insight into which aspects of textual input, such as stylistic instructions or descriptive terms, have the most significant impact on the final output.
The first step utilizes a text generation model that is trained on input text. The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output. The text generation model is expected to produce responses similar to the original response, varying only based on the critical components of the input texts [117].
Different text permutations are utilized from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into separate words. The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [118].
While the text generation model faces challenges in tasks such as counting objects and spatial reasoning, these issues can be mitigated by carefully selecting perturbation texts and restricting the input domain for perturbations [119]. The text generation models enable the proposed method to create responses that differ from the original text solely in terms of keywords related to the perturbation texts.
We use the Wasserstein distance to quantify these differences, comparing the distributions in both text embeddings [120, 112]. This metric is employed because it provides a more comprehensive measure of the differences between text embeddings compared to other ECDF-based distances, such as Kolmogorov–Smirnov or Cramér–von Mises. Although these methods focus on specific aspects, the Wasserstein distance captures both broad and subtle changes in the data [120, 113].
Finally, by applying linear regression, the space of input prompts is linked to the output space of the text generation model. As a result, the model’s behaviour can be interpreted by examining the regression coefficients [121, 116].
4.1 Text Generating and Perturbed Prompts
To begin, we generate variations of the original text prompt, referred to as “perturbed prompts,” to analyze how subtle changes in the input affect the output of a text generation model [122, 123]. The original prompt $p_{\text{org}}$ is first tokenized into individual words. From these tokens, multiple perturbed prompts $\{p_{\text{pert}}^{i}\}_{i=1}^{N}$ are constructed by selectively including or excluding specific words. Each perturbed prompt is then passed to the text generation model, which produces a corresponding output. The output generated from the original prompt, $t_{\text{org}}^{\prime}$, serves as the reference (or baseline) for comparison.
For the original text prompt $p_{\text{org}}$, the text generation model $\phi_{\text{gen}}$ produces the output:
$t_{\text{org}}^{\prime}=\phi_{\text{gen}}(p_{\text{org}})$ (11)
Similarly, for each perturbed prompt $p_{\text{pert}}^{i}\in S_{\text{pert}}$, the model generates:
$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in\{1,\dots,N\}$ (12)
Here, $\phi_{\text{gen}}$ is the text generation function, and $t_{\text{pert}}^{\prime i}$ denotes the generated output for the $i$-th perturbation. These outputs are used to measure how individual word-level changes in the prompt influence the model’s behavior [63, 118].
4.2 Creating the Interpretable Space
To represent the output space of the text generation model in a one-dimensional similarity space, we compute the Wasserstein distance between each generated text from a perturbed prompt and the output generated from the original prompt [112, 113]. Throughout this process, the input structure remains unchanged, and only the perturbed words differ across inputs.
To capture the semantic similarity between generated outputs, we use Word Mover’s Distance (WMD) to embed the texts into a feature space [115]. These embeddings enable the focus to remain on high-level semantic content, thereby mitigating lexical noise and enhancing the reliability of the computed distances. The resulting scalar distances are used as the dependent variable in the subsequent regression model.
The Wasserstein distance between each perturbed output $t_{\text{pert}}^{\prime i}$ and the original output $t_{\text{org}}^{\prime}$ is computed as:
$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{P}\right)^{\frac{1}{P}},\quad\forall i\in\{1,\dots,N\}$ (13)
In this equation:
  • $\delta_{\text{WMD}}(\cdot)$ denotes the semantic embedding of a generated text,
  • $n$ is the dimensionality of the embedding space,
  • $P$ specifies the norm order used in the Wasserstein metric,
  • and $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance computed for the $i$-th perturbation.
Using p-values is crucial for validating the significance of the observed Wasserstein distances (WD) between texts generated under different prompt perturbations. By calculating a p-value, we test whether the observed WD could have occurred by chance, determining if the perturbation-induced changes are meaningful [124]. To compute this p-value, we use the Bootstrap algorithm [125, 126], which estimates the probability of observing a WD as large as the observed one. This involves generating distributions of WDs through repeated sampling and comparing each sampled WD with the observed value. If a small proportion of sampled WDs exceed the observed WD, the resulting p-value confirms that the differences are statistically significant, supporting the validity of the perturbation effects.
The Wasserstein-$P$ metric enables adjustment of the sensitivity of the Wasserstein Distance (WD) to different types of text variations by tuning the norm order $P$. Smaller values of $P$ emphasize fine-grained, local differences between distributions, while larger values capture broader, more global shifts [113]. By experimenting with various values of $P$, we can select the one that best captures the semantic differences introduced by prompt perturbations, thereby improving both the reliability and interpretability of the analysis [113].
To assess the statistical significance of the computed Wasserstein distances, we apply hypothesis testing using p-values. Distances associated with high p-values, indicating non-significant variation, are excluded from further consideration [124]. As detailed in Algorithm 1, the Bootstrap algorithm [127, 108] is used to estimate these p-values for each observed WD.
For example, let X and Y denote two sets of text-based embeddings. The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y. In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed. If boostWD exceeds the original observed WD, a counter (bigger) is incremented. After all iterations, the p-value (pVal) is estimated as the proportion of iterations where boostWD exceeds the observed WD, as outlined in Algorithm 1.
Input: Two sets of text-based distances: $X$ and $Y$
Output: Observed Wasserstein Distance (WD) and corresponding p-value (pVal)
1
[…]
6 $n\leftarrow\texttt{LX}+\texttt{LY}$;
7 $\texttt{bigger}\leftarrow 0$;
8
9for $i\leftarrow 1$ to MaxItr do
10       Draw a random sample $e$ of size LX from $XY$;
11       Draw a random sample $f$ of size LY from $XY$;
[…]
14             $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
15
16       end if
17
18 end for
[…]
Algorithm 1 Bootstrap Algorithm for Wasserstein Distance (WD) P-Value Calculation
Figure 5: gSMILE flowchart for Large Language Models
When generating perturbation texts from the original prompt, the key terms in the prompt are essential for maintaining semantic similarity between perturbed prompts and the source. To account for this, we define a similarity metric between each perturbed prompt and the original, which is then used as a sample weight in the interpretable model. This similarity is quantified using Word Mover’s Distance (WMD), which calculates the minimum cost to transport the embedded words of one document to match another [115].
The WMD between the original prompt $p_{\text{org}}$ and a perturbed prompt $p_{\text{pert}}^{i}$ is defined as:
$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\cdot d(w_{k},w_{l})$ (14)
subject to $\sum_{l=1}^{m}T_{kl}=d_{k},\quad\forall k=1,\dots,n$
$\sum_{k=1}^{n}T_{kl}=d^{\prime}_{l},\quad\forall l=1,\dots,m$
$T_{kl}\geq 0$
Here:
  • $w_{k}$ and $w_{l}$ denote the $k$-th and $l$-th words from the original and perturbed prompts respectively,
  • $d(w_{k},w_{l})$ is the distance (typically cosine or Euclidean) between the embeddings of $w_{k}$ and $w_{l}$,
  • $T_{kl}$ is the flow variable indicating how much of $w_{k}$’s weight is transported to $w_{l}$,
  • $d_{k}$ and $d^{\prime}_{l}$ represent the normalized bag-of-words weights for the original and perturbed texts.
To use this distance as a weighting factor in the regression model, we normalize it using a Gaussian kernel:
$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})=\exp\left(-\left(\frac{\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma}\right)^{2}\right),\quad\forall i\in\{1,\dots,N\}$ (15)
Here, $\sigma$ is a kernel width parameter that controls sensitivity to dissimilar inputs. The resulting similarity score $\pi_{i}$ falls in the range $[0,1]$, assigning higher weights to perturbed prompts that are semantically closer to the original.
4.3 Developing the Interpretable Surrogate Model
To interpret the influence of individual words on text generation, we map the output space of the generation model to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [128]. This enables the training of a local surrogate model that relates changes in input structure to variations in output behavior.
In this weighted linear regression framework, the perturbation vectors (i.e., binary or one-hot encodings of included or excluded tokens) serve as independent variables. The response variable is the Wasserstein distance between the output of a perturbed prompt and that of the original prompt. Each perturbation is assigned a weight based on its semantic similarity to the original prompt [114, 129].
The loss function minimized during training is a weighted squared error:
$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\cdot\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$ (16)
Here:
  • $\phi_{\text{gen}}$ is the text generation model,
  • $\delta_{\text{WMD}}$ denotes the function used to embed generated texts,
  • $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance between the generated outputs,
  • $\pi_{i}$ is the similarity-based kernel weight for the $i$-th perturbation,
  • and $\hat{f}_{t}(p_{\text{pert}}^{i})$ is the prediction from the linear surrogate model.
The surrogate model $\hat{f}_{t}$ is trained on the one-hot vector representation of the perturbation prompts $p_{\text{pert}}^{i,01}$, using the computed Wasserstein distances as targets [114].
4.4 Evaluation Metrics
In our investigation, we adopt a suite of evaluation metrics inspired by foundational work provided by Google [130], emphasising the multifaceted nature of assessing explainable models. This work highlights the significance of metrics such as accuracy, stability, fidelity and consistency as essential tools for a rigorous evaluation of model behaviour, particularly when comparing explainable models to traditional black-box models. The adoption of these metrics provides a structured methodology to dissect and understand model reliability in a more holistic manner [130].
Figure 6:
4.4.1 Accuracy
Accuracy measures how well the model’s output aligns with the expected results (ground truth). Specifically, we compare the model’s attribution scores (attention to text elements) against ground truth labels that identify the most relevant text elements for specific generated text segments [131].
To quantify this, we use the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) [132]. AUC reflects the model’s ability to rank relevant elements (from the ground truth) higher than irrelevant ones:
  • AUC $\sim$1: The model effectively distinguishes relevant elements, closely matching human-identified ground truth.
  • AUC $\sim$0.5: The model’s ranking is random, showing poor alignment with the ground truth.
For example, Fig. 6 shows that the ground truth identifies the words “meaning” and “life” as the most relevant text elements for generating the corresponding text output. The metric AttAUC evaluates the model’s alignment with the ground truth, where:
  • AttAUC = 1.0: Perfect alignment.
  • Lower AttAUC (0.8): Less accurate [133].
The heatmaps produced by the model visually represent these attention scores, as shown in Fig. 6, where darker red shades indicate higher attribution to specific text elements. An accurate model should correctly assign higher attention to the text elements that most influence the generated text output, improving the interpretability and reliability of the model [131].
4.4.2 Stability
Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [134]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. Stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [10].
In the example text, the attention scores are visualised for two slightly different text prompts. The heatmaps show how the model assigns attention to each word. A stable model, as demonstrated in Fig. 6, assigns similar importance to the words make and rainy in both prompts, even though the token  was added in one case. This consistency is critical for ensuring the model behaves predictably across minor input variations [136]. Without such stability, the explanations generated by the model could lead to confusion or misinterpretation, especially in high-stakes scenarios such as medical diagnosis or financial decision-making [63].
We use the Jaccard index to quantify stability, which measures the similarity between two sets of model predictions. For two sets $A$ and $B$, the Jaccard index is calculated as follows in Eq. 17:
$\text{Jaccard}(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (17)
This metric reflects the proportion of shared elements between the sets, where values closer to 1 indicate greater stability and similarity in explanations [134]. A higher Jaccard index confirms the model’s ability to yield stable outputs, thus providing consistent and reliable interpretations [4]. By ensuring stability, we not only enhance the robustness of the model but also its usability and credibility in real-world applications [61].
4.4.3 Consistency
In evaluating the consistency of our model, we assess whether it produces stable and reliable outputs when the same input is provided multiple times. Consistency ensures that the model behaves predictably, reinforcing trust in its outputs and usability across diverse scenarios [63].
As shown in Fig. 6, when the input phrase “what is the meaning of life?” was fed into the model multiple times, the output remained stable across iterations. This consistency is crucial for applications where repeatable and reliable results are required. In our case, the model consistently assigns higher weights to the words “meaning” and “life”, demonstrating that it understands the importance of these words in generating the desired output [4]. Such behaviour underscores the model’s ability to focus on the semantic core of the input, a trait essential for maintaining coherence in practical use cases [10].
In practical applications—such as automated content generation, interactive systems, or scenario simulations—users depend on models to produce results that do not fluctuate unexpectedly. Any inconsistencies could undermine user trust and compromise the system’s overall effectiveness. Our model enhances user confidence by demonstrating consistent outputs and establishes a foundation for further optimisation and integration into more complex systems. This ability to maintain consistency across iterations directly contributes to the robustness and scalability of the model, making it well-suited for deployment in dynamic real-world environments [61].
4.4.4 Fidelity
Fidelity is a metric used to evaluate how closely an explainable model replicates the behavior of a black-box model. Specifically, the predictions of the black-box model denoted as $f(X_{i})$, are compared to those of the explainable model, $g(X_{i})$, for each input perturbation $i$. A variety of accuracy-based and error-based metrics are employed to assess this alignment, thereby quantifying how well the explainable model approximates the responses of the black-box model.
Let $N_{p}$ denote the total number of perturbation samples (i.e., data points). One of the most commonly used accuracy metrics is the coefficient of determination $R^{2}$, which measures how well the explainable model’s predictions correlate with those of the black-box model [135]. It is defined as:
$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f})^{2}}$ (18)
where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{i})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
In cases where data samples have different levels of importance, a weighted variant is used. The weighted coefficient of determination $R_{w}^{2}$ is given by [136, 137]:
$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}_{w})^{2}}$ (19)
Here, $\bar{f}_{w}$ is the weighted mean of the black-box predictions. These accuracy-based metrics ($R^{2}$ and $R_{w}^{2}$) provide insight into how closely the explainable model aligns with the black-box model.
To account for both sample size and model complexity (i.e., number of features used), the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ is used [136]:
$\hat{R}_{w}^{2}=1-(1-R_{w}^{2})\left[\frac{N_{p}-1}{N_{p}-N_{s}-1}\right]$ (20)
Where $N_{s}$ is the number of features in the interpretable model.
In addition to correlation-based metrics, error-based metrics are widely used to quantify differences in prediction values. The classical unweighted forms Mean Squared Error (MSE) and Mean Absolute Error (MAE) are:
$\text{MSE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (21)
$\text{MAE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (22)
For situations requiring sample-level importance, the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE) are defined as [138]:
$\text{WMSE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}w_{i}}$ (23)
$\text{WMAE}=\frac{\sum_{i=1}^{N_{p}}w_{i}\cdot|f(X_{i})-g(X_{i})|}{\sum_{i=1}^{N_{p}}w_{i}}$ (24)
Here, $w_{i}$ denotes the sample weight associated with the $i$-th perturbation.
Moreover, the mean $L_{1}$ and mean $L_{2}$ norms are also applied to evaluate absolute and squared differences respectively [139, 136]:
$L_{1}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (25)
$L_{2}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (26)
These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$—together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
We compute fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model. This includes analyzing how different input perturbations affect fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behavior.
5 Experimental Results
This section evaluates the proposed method’s ability to enhance explainability in instruction-based text generation. We analyse how the structure of textual prompts affects interpretability, how variations in input text influence the clarity of explanations, and how the proposed solution performs across diverse scenarios.
Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI GPT [1], LLaMA [19], and Claude-AI [140] to ensure a comprehensive assessment.
5.1 Qualitative Results
The qualitative evaluation demonstrates the proposed framework’s interpretability using scenario-based testing and visualisations. We designed diverse testing scenarios by varying textual prompts, including changes in sentence complexity, rephrasing, and the inclusion of domain-specific terminology. These scenarios examine how the model interprets and generates outputs based on nuanced variations in textual inputs.
5.1.1 Input Prompt Explainability
The proposed framework applies scenario-based testing across multiple prompt variations to evaluate the interpretability of the model’s behaviour regarding different inputs. These variations include rephrased questions, descriptive modifiers, and changes in word choice. The system generates a text heatmap highlighting individual words’ contribution to the model’s output for each input, with colours indicating the importance scores.
Figure 7 shows how varying the input prompt’s phrasing or focus leads to different word importance patterns, demonstrating the model’s sensitivity to linguistic changes.
Figure 7: Text heatmaps showing word-level importance for various input prompts. The colour intensity represents the influence of each word on the model’s output.
5.1.2 Reasoning Path and Token Attribution
Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response. Figure 8 presents an example where the gSMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram. This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models. As shown in Figure 8, gSMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [141].
Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the gSMILE framework. The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates gSMILE’s token-level importance heatmap for the same prompt. [141]
5.1.3 Word-Level Attribution Analysis on MMLU Using gSMILE
To evaluate the interpretability of ChatGPT using the gSMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [142]. This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities. For this study, we accessed a pre-formatted version of the dataset available on Kaggle [143] to facilitate prompt construction and experimentation.
Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision. Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter. These visualisations demonstrate gSMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
Figure 9: Attribution heatmaps generated by gSMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
5.1.4 Optimised Instructions from Google Research
To further explore how instruction phrasing influences interpretability, we applied gSMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [144]. These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
We first applied gSMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem. ” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, gSMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using gSMILE with GPT-3.5.
Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using gSMILE with GPT-4.
5.2 Quantitative Results
In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to assess our explainability method across different text generation models and different scenarios.
5.2.1 Accuracy
To evaluate the performance of instruction-based text generation models, we tested ten different prompts and their variations through 64 perturbations. Heatmaps were generated to display the weights of each keyword in the prompts across various models, including OpenAI GPT, LLAMA, and Claude-AI. A ground truth was defined, assigning weights of 1 to critical keywords in the prompts and 0 to less relevant or auxiliary words. These ground truth values were compared with the extracted heatmaps to assess the interpretability and accuracy of each model.
To quantify accuracy, we employed multiple metrics, including Attention Accuracy (ATT ACC), F1-Score for Attention (ATT F1), and Area Under the Receiver Operating Characteristic Curve for Attention (ATT AUROC). The results were averaged across the ten different prompts, and each model’s performance is summarised in Table 1, showcasing their effectiveness based on these metrics.
Among the evaluated models, Claude-AI demonstrated the strongest performance, achieving the highest scores in all three metrics: 0.82 in ATT ACC, 0.67 in ATT F1, and 0.88 in ATT AUROC. These results indicate Claude-AI’s superior ability to consistently focus on essential input tokens across varying prompts. In contrast, LLAMA achieved the second-highest accuracy but showed a notably lower F1-score (0.40), suggesting potential inconsistency in identifying key tokens. OpenAI GPT yielded balanced results but lagged slightly behind Claude-AI in all metrics, highlighting room for improvement in its attention fidelity. Ll, Claude-AI provides more reliable and semantically focused attention compared to its counterparts.
Table 1: Performance metrics (ATT ACC, ATT F1, and ATT AUROC) for various
models, evaluated under two conditions: different prompts and different images.
Model name ATT ACC ATT F1 ATT AUROC
OpenAI-GPT 0.70 0.59 0.84
LLaMA 0.76 0.40 0.84
Claude-AI 0.82 0.67 0.88
5.2.2 Stability
To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap. For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word. To evaluate stability, we added only the token  at the end of the text, extracted the coefficients again, and computed the Jaccard index between the coefficients before and after the perturbation. The averaged Jaccard indices across all ten prompts for each model are reported in Table 2. Higher scores indicate greater stability under minor input changes.
Table 2: Stability scores across models based on average Jaccard index over ten prompts
Model name Jaccard Index
OpenAI-GPT 0.62
LLaMA 0.45
Claude-AI 0.44
5.2.3 Consistency
To compute consistency, we ran the same code with 64 perturbations for each prompt over 10 iterations, ensuring that the outputs remained consistent across these repetitions for text generation models such as OpenAI GPT, LLAMA, and Claude-AI, as shown in Table 3. We computed variance and standard deviation metrics for each word coefficient. This step is essential to confirm that the model’s predictions are not random or heavily dependent on initialisation factors but are instead deterministic and robust.
Among all models, OpenAI-GPT exhibited the highest consistency, achieving a variance of 0.0000 and a standard deviation of 0.0046. This indicates that its predictions are the most stable across repeated runs, confirming its robustness and reliability in generating interpretable outputs.
Table 3: Consistency metrics for different models on the prompt
“What is the meaning of life?”
Model name Variance Standard Deviation
OpenAI-GPT 0.0000 0.0046
LLaMA 0.0048 0.0663
Claude-AI 0.0011 0.0331
5.2.4 Fidelity for Different Instruct image editing diffusion based Models (IED)
[…]
Table 4: Fidelity metrics for different models on the prompt
“What is the meaning of life?”
[…]
OpenAI-GPT 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
LLaMA 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
Claude-AI 0.0265 0.6967 0.1251 0.1981 0.0708 0.6240
From the results, we observe that Claude-AI achieves the lowest WMSE, WMAE, and mean-L1, indicating superior performance in minimizing error. On the other hand, OpenAI-GPT performs best in terms of $R^{2}_{\omega}$, mean-L2, and adjusted $R^{2}_{\hat{\omega}}$, demonstrating stronger alignment and generalization of the surrogate model. LLaMA ranks in the middle across most metrics but does not lead in any. This suggests that different models excel in different fidelity dimensions, highlighting the importance of multi-metric evaluation when interpreting black-box model behavior.
5.2.5 Fidelity Across Different number of Text Perturbations
Tables 56, and 7 present the results of fidelity computations for varying numbers of perturbations across the OpenAI GPT, LLAMA, and Claude-AI models, respectively. These results are derived using a weighted linear regression as the surrogate model and the Wasserstein distance for computing distances. The evaluation metrics include weighted mean squared error (WMSE), weighted mean absolute error (WMAE), coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$), and average loss metrics such as mean-L1 and mean-L2.
Table 5: Performance metrics for different numbers of perturbations in OpenAI GPT
[…]
32 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
64 0.0617 0.3406 0.1695 0.1756 0.0526 0.2712
128 0.0329 0.4468 0.1385 0.1519 0.0378 0.4193
256 0.0116 0.4276 0.0675 0.0872 0.0163 0.4138
Table 6: Performance metrics for different numbers of perturbations in LLaMA
[…]
32 0.0368 0.7068 0.1617 0.2387 0.0805 0.6364
64 0.0240 0.4844 0.1284 0.1433 0.0382 0.4301
128 0.0085 0.2430 0.0638 0.0612 0.0096 0.2055
256 0.0058 0.2387 0.0605 0.0667 0.0118 0.2203
Table 7: Performance metrics for different numbers of perturbations in Claude-AI
[…]
32 0.0368 0.7209 0.1495 0.2262 0.0733 0.6539
64 0.0115 0.3708 0.0564 0.1036 0.0370 0.3046
128 0.0089 0.1506 0.0535 0.0920 0.0192 0.1085
256 0.0191 0.2670 0.0902 0.0873 0.0163 0.2494
For OpenAI GPT and LLaMA, increasing the number of perturbations generally improves fidelity across most metrics, with 256 perturbations achieving the best WMSE, WMAE, and loss values. Interestingly, for Claude-AI, 64 perturbations yield the lowest WMSE and WMAE, while 256 achieves the best loss (mean-L2 and mean-L1). However, the highest $R^{2}$ scores for all three models occur at 32 perturbations, suggesting that beyond a certain point, more perturbations may dilute interpretability despite improving local fit. Therefore, there is a trade-off between the number of perturbations and the reliability of the surrogate model, depending on the metric of interest.
5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
Fidelity is also computed by comparing various distance metrics for text generation models such as OpenAI GPT using text-to-text comparisons. The surrogate models used are Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME), as shown in Table 8. We explore different combinations of distance metrics, including Cosine similarity and Wasserstein Distance (WD). Fidelity is measured using WMSE, WMAE, mean-$L_{1}$, mean-$L_{2}$, and coefficient of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$).
Table 8: Fidelity results for different distance measures with 30 perturbations
WLR Fidelity Metrics
[…]
Cosine Cosine 0.0172 0.3151 0.0659 0.1277 0.0412 0.1508
Cosine WD 0.0216 0.4197 0.0899 0.1332 0.0385 0.2805
WD WD 0.0388 0.7104 0.1731 0.2035 0.0609 0.6409
WD Cosine 0.0048 0.4026 0.0329 0.0871 0.0296 0.2593
WD+C WD+C 0.0349 0.5349 0.1589 0.3050 0.1468 0.4233
BayLIME
Cosine Cosine 0.0200 0.2285 0.0499 0.1067 0.0464 0.0434
Cosine WD 0.0241 0.3572 0.0736 0.1281 0.0430 0.2029
WD WD 0.0118 0.5837 0.0629 0.1170 0.0361 0.4838
WD Cosine 0.0058 0.5273 0.0459 0.0978 0.0304 0.4139
WD+C WD+C 0.0048 0.4227 0.0284 0.0690 0.0213 0.2841
Our Proposed Method: gSMILE.
Based on the results in Table 8, the combination where Wasserstein Distance (WD) is used for both input similarity and output shift consistently yields the highest values of the weighted coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$). Specifically, underweighted linear regression, the WD–WD configuration achieves $R^{2}_{\omega}=\textbf{0.7104}$ and $R^{2}_{\hat{\omega}}=\textbf{0.6409}$, which significantly outperform other configurations.
Motivated by these findings, we introduce our method, called gSMILE (generalized SMILE), which utilizes the Wasserstein distance for both input and output comparisons. This choice enhances the surrogate model’s fidelity and stability in capturing local shifts in model behavior.
We use gSMILE throughout subsequent sections as the foundation for explainability evaluations on large language models (LLMs). These include assessments of fidelity, consistency, and stability, ensuring the proposed approach is both theoretically grounded and empirically adequate.
5.2.7 Computation Complexity
We evaluate the computational efficiency of our proposed method, gSMILE, using three instruction-based text generation models: OpenAI GPT, LLaMA, and Claude-AI. Although these models differ in scale and architecture, the objective here is not to compare the models themselves but rather to assess the performance of different explainability methods applied to each model under identical hardware and runtime conditions. All evaluations are conducted with 60 perturbations per input prompt.
We compare three explainability approaches—LIME, gSMILE, and Bay-LIME—based on their execution time. As reported in Table 9, gSMILE consistently requires the highest runtime across all model frameworks. This is mainly due to its reliance on the Wasserstein distance to compute similarity between output distributions, which involves solving high-dimensional optimization problems with time complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [145]. By contrast, LIME and Bay-LIME utilize cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
Despite the higher computational cost, gSMILE delivers significantly better fidelity, stability, and robustness to adversarial perturbations—essential properties for trustworthy explanations. These advantages are supported by results in Tables 8 and 2 and align with findings from prior work on SMILE [11], which demonstrated superior robustness against adversarial input variations.
These findings suggest that gSMILE’s computational overhead is justified by the quality of explanations it provides, making it highly suitable for applications where interpretability, consistency, and resistance to manipulation are paramount.
Table 9: Execution times (in seconds) for each framework and explainability method with 60 perturbations
Framework LIME gSMILE Bay-LIME
OpenAI-GPT 156.82 170.70 161.43
LLaMA 113.00 118.99 116.40
Claude-AI 355.09 372.95 349.88
5.2.8 Token-Level Attention for Different Sentence Structures
To investigate how sentence structure affects token-level attention in large language models (LLMs), we analyzed the attention scores assigned to the tokens “meaning” and “life” across ten semantically similar but syntactically varied prompts. These prompts ranged from questions (e.g., “What is the meaning of life?”), commands (e.g., “Please give me the meaning of life”), to declarative formulations (e.g., “You must explain the meaning of life”).
For each sentence, we extracted the attention values directly from the final layer of the LLM. We focused on the attention allocated to the tokens “meaning” and “life.” The results were visualized using a box plot (see Figure 12), where each point represents a unique sentence. Full-sentence prompts are viewable through hover text, allowing qualitative comparison of sentence structure with attention distribution.
The analysis reveals noticeable variability in how these tokens are weighted depending on syntactic form. For instance, in imperative constructions, attention to “life” tends to dominate, whereas in interrogative or abstract prompts, “meaning” often receives comparable or higher attention. This suggests that sentence formulation can subtly steer the internal focus of the model, even when the semantic content remains consistent.
These findings underscore the sensitivity of LLMs to surface-level structure in shaping token-level interpretability, offering insight into both model behaviour and prompt engineering strategies.
Figure 12: Box plot showing attention scores for the tokens “meaning” and “life” across ten prompts with varying sentence structures. Each point represents the attention score assigned to the token in a specific sentence instance.
6 Limitations
While gSMILE provides a flexible and model-agnostic framework for interpreting large language models (LLMs), its application introduces several limitations due to the use of perturbation-based sampling and LLMS’ black-box nature.
1) Perturbing prompts at the token level can lead to disproportionately large changes in LLM outputs, particularly in English, even when the semantic meaning remains largely intact. This sensitivity challenges the assumption of local smoothness required for regression-based attribution and can reduce explanation stability.
2) Some LLMs exhibit hallucination, producing fluent but factually incorrect or irrelevant outputs, especially when presented with ambiguous or syntactically degraded perturbations. This introduces noise into the attribution signal and affects explanation reliability.
3) The computational cost of generating and processing many perturbations (e.g., 1024 per input) is significant. Commercial APIs can lead to high inference costs and latency. For local models, substantial GPU resources are required, limiting practical scalability.
4) Some models, particularly research-stage or open-source LLMs, do not provide API access and require manual local deployment.
7 Conclusion
Interpretability remains a cornerstone of responsible AI, particularly in the context of large language models (LLMs) and instruction-based text generation. As AI systems become increasingly embedded in domains such as business, education, healthcare, and public services, it becomes essential for users to understand how these models generate outputs from their prompts. Without such insight, trust in AI systems diminishes, limiting adoption in real-world workflows. This underscores the need for transparent and interpretable solutions to demystify the behaviour of these black-box models.
Interpretability is not only a research priority but is likely to become a legal and operational requirement in alignment with emerging AI regulations and ethical standards. Organisations that fail to address interpretability may encounter regulatory hurdles and reduced user acceptance.
This work introduced a novel interpretability framework tailored for instruction-based text generation models. Leveraging gSMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output. Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
Whereas most existing interpretability methods highlight important tokens in the model’s output [63, 64], gSMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma. The resulting heatmaps and attribution scores provide an intuitive understanding of model behavior, empowering users to control and refine prompt design better. gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.
This work lays the groundwork for more transparent and trustworthy natural language systems. Future directions include extending gSMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
8 Future Works
  1. 1.
    Extension to Mathematical and Equation-Based Text Generation: Future work will explore the application of the proposed interpretability framework to models focused on symbolic reasoning and math-related tasks, such as equation solving, theorem generation, and formal derivation. Instruction-based mathematical models like Minerva and MathGLM require high precision in prompt design, and understanding how prompt components affect symbolic output will be essential in educational and scientific settings.
  2. 2.
    Application to Paraphrase Generation Models: The methodology can be extended to paraphrasing models, where semantic preservation and variation are central. Analysing how individual words influence paraphrased outputs’ style, tone, or fidelity can support better user control in writing assistance, content rephrasing, and translation tools.
  3. 3.
    Adaptation to Multi-Turn and Context-Aware Generation: Another direction is to adapt the framework for multi-turn and context-aware systems, including dialogue agents. This would involve tracking how prompt changes impact responses’ consistency, coherence, and relevance across conversational sequences.
  4. 4.
    Cross-Task Generalisation: While this work focuses on instruction-based text generation, the interpretability strategy can be generalised to broader NLP tasks, including summarisation, classification, and question answering. Evaluating how different model families respond to prompt perturbations can uncover model-specific behaviours and improve transparency.
  5. 5.
    Benchmarking Interpretability and Prompt Sensitivity: We plan to contribute to standardised evaluation protocols for prompt sensitivity, attribution robustness, and local fidelity. These benchmarks will support consistent assessment across models and help identify best practices for interpretable generation.
  6. 6.
    Optimisation of Computational Efficiency: To address the computational demands of gSMILE, future work will investigate dimensionality reduction techniques and approximation strategies. These approaches aim to reduce execution time while preserving the integrity and interpretability of the model’s outputs.
Data Availability
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the gSMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.